For its first year of operations, Tringali Corporation's reconciliation of pretax accounting income to taxable income is as follows:

Pretax accounting income$ 230,000
Permanent difference(15,200)
 214,800
Temporary difference-depreciation(19,000)
Taxable income$ 195,800

Tringali's tax rate is 25%. Assume that no estimated taxes have been paid.

What should Tringali report as its income tax expense for its first year of operations?
Answer #1

step 1) 195,800*25%= 48,950

step 2) 19,000*25%= 4,750                          

48,950+4,750= 53,700 Answer


answered by: anonymous